OpenStudy (anonymous):

Solve: 0 = 4x^2 − 8x − 5.

Parth (parthkohli):

\[4x^2 - 10x + 2x - 5 = 0\]

OpenStudy (gorv):

(2x-5)(2x-1)=0 x=5/2 and x=1/2

OpenStudy (gorv):

x=-1/2

OpenStudy (gorv):

not 1/2

OpenStudy (mayankdevnani):